Boat building, iron foundries, and amusement parks and arcades are three of the industries with the highest rates of occupational injuries. The graph shows the number of nonfatal injuries per 1000 full-time workers in 2004 in these three industries. In a typical group of 1000 workers from each of these industries, there were 30 more injuries in iron foundries (I) than in amusement parks and arcades (A), and there were 12 more injuries in amusement parks and arcades than in boat building (B). Among these 3000 workers, there were 387 nonfatal occupational injuries. How many of the injuries took place in each industry?
Answer by macston(5194)

A=injuries in amusement parks; B=injuries in boat building=A-12: I=injuries in iron foundries=A+30
A+B+I=387 Substitute for B and I
A+(A-12)+(A+30)=387
3A+18=387 Subtract 18 from each side.
3A=369 Divide each side by 3.
A=123 ANSWER 1: There were 123 injuries in Amusement Parks and Arcades.
B=A-12=123-12=111 ANSWER 2: There were 111 injuries in Boat Building.
I=A+30=123+30=153 ANSWER 3: There were 153 injuries in Iron Foundries.